Grade upto college level Physical Chemistry

Naturally occurring consist of two isotopes whose atomic weights are 10.01 and 11.01. The atomic weight of natural boron is 10.81. Calculate the percentage of each isotope in natural boron.

Shane Macguire
12 Years agoGrade upto college level

1 Answer

Deepak Patra
12 Years ago
Let the % of isotope with At. wt. 10.01 = x
∴ % of isotope with At. wt. 11.01 = (100 - x)
At. wt. of boron = x *10.01 + (100 - x) *11.01/100
=> 10.81 = x *10.01 + (100 - x) *11.01/100
∴ x = 20
Hence % of isotope with At. wt. 10.01 = 20%
∴ % of isotope with At. wt. 11.01 = 100 – 20 = 80%.